# Xpikeformer: Hybrid Analog-Digital Hardware Acceleration for Spiking Transformers

### v1

@@ IV-C Dataflow
The AIMC engine serially drives input data through each forward propagation or fully connected layer. Each AIMC cycle, except for the first and last, includes two feedforward layers and one fully connected layer to project embeddings into the $QKV$ feature space. In each layer, data is input in a “token-wise event-driven” manner: the spike encoding of the same token embedding is input sequentially over $T$ time steps, followed by cycling through $N$ tokens. This approach allows the membrane potential stored in the LIF unit’s register to accumulate directly and reset upon token switching. Looping $N$ tokens for each time step $t$ would require storing and reading back membrane potentials, leading to unnecessary data transmission overhead.
@@
The SSA engine operates multiple SSA tiles in parallel, with each tile handling one head of the multi-head attention mechanism. As previously discussed, $Q$, $K$, and $V$ data in SSA tiles are transmitted in a “matrix-wise event-driven” manner. The three matrices $Q^{t}$, $K^{t}$, and $V^{t}$ are streamed to the SSA tile column by column over $d_{K}$ clock cycles. The subsequent matrices $Q^{t+1}$, $K^{t+1}$, and $V^{k+1}$ are transmitted in the next $d_{K}$ clock cycles. When pipelining in time steps, the computational delay of an SSA tile (from first input to first output) is approximately $d_{K}$, ensuring high computational efficiency of the SSA engine.
@@ V-A Integrated Framework for Hybrid Training
We adopt a two-stage training procedure – at the initial training stage, conventional training (CT) is used in an ideal full-precision digital environment using SpikingJelly. After obtaining the trained parameters, AIHWKit is used to conduct hardware-aware training (HWAT) to fine-tune the parameters. This is achieved by injecting noise in the forward propagation while keeping the backward propagation ideal. This procedure adapts the network parameters to realistic PCM-based hardware imperfections.
During training, the forward propagation outputs are averaged over time steps to calculate the loss, with the loss function defined similarly to that in training ANN transformers. The Adam optimizer is employed to train the Xpikeformer network [49].
@@ V-B Global Drift Compensation
To address the conductance drift of the PCM devices over time in the AIMC engine, we implement a global drift compensation (GDC) method introduced in [50]. During the calibration phase, we periodically measure the total current of several columns in each SA programmed to known conductance values. This allows Xpikeformer to track global shifts in conductance. During inference, the outputs are scaled by a factor derived from these calibration measurements. The calibration can be performed while the PCM array is idle, with the summation of currents implemented within the control unit of the system.
@@ VI-A Tasks Description
The inference accuracy of Xpikeformer is validated in hardware-simulated environments using AIHWKit. We assess the encoder-only and decoder-only architectures separately through two respective tasks.
Task 1 (Image Classification): To evaluate the accuracy performance of the encoder-only Transformer, we select an image classification task as it requires the model’s ability to extract and process relevant features from the entire input image in parallel. We utilized an SNN-based vision transformer (ViT). We selected the CIFAR-10 dataset, which comprises 10 different classes of objects.
Figure 7: A conceptual illustration of in-context learning (ICL) using a decoder-only transformer.
Task 2 (In-context Learning for Wireless Communication): Decoder-only transformers are often used for tasks involving temporal sequential dependencies. A notable ability of these models is in-context learning (ICL). As illustrated in Fig. 7, the model is provided with a sequence of query-answer pairs that follow a consistent pattern, even though the model has not been explicitly trained on this pattern. The model is then expected to predict the answer for a new query by attending to the preceding query-answer pairs within the input sequence. This ability has been demonstrated to enable wireless receivers to correctly classify symbols after they have undergone unknown wireless channels and been contaminated by nonlinearities [3, 4]. We adopt the symbol detection task as formulated in [27] and evaluate accuracy performance using the bit error rate (BER) metric (a lower BER indicates higher accuracy). We fix the number of input query-answer pairs to 18. An SNN-based GPT-2-like decoder-only transformer is employed, with size parameters detailed in Table III.
TABLE III: Model Configuration Parameters for Task 1 and Task 2
Parameter Task 1 Task 2
Token dimension 48 4
Embedding dimension 512 256
Hidden dimension 2048 1024
# layers 6 4
# heads 8 8
# classes 10 16
Attention mask Casual mask Diagonal mask
VI-B Accuracy Results
Due to the noise, variability, and non-linearities of analog computations, the inference accuracy of Xpikeformer is inherently lower than that of full-precision digital computations. For Task 1, we train the SNN-based ViT on the CIFAR-10 training set from scratch, using the HWAT strategy as introduced in Section V-A. We test the inference accuracy on the evaluation dataset using the simulation platform as described in Section V-B. For Task 2, the decoder-only model is trained on a training dataset that is generated from a set of $8192$ wireless channels and tested on a dataset generated from another $1024$ wireless channel distinct from the training dataset. We compare the Xpikeformer results with two baseline implementations: (i) ANN (Digital) – A traditional ANN transformer trained and tested on fully digital devices; (ii) SNN (Digital) – an SNN transformer trained and tested on fully digital devices, with its attention implemented by SSA. The parameters of both baselines are INT8-quantized.
For Task 1, we evaluated image classification accuracy over different spike encoding lengths, $T$. Fig. 8 presents the accuracy of SNN (Digital) and Xpikeformer across varying spike encoding lengths, compared to the baseline digital ANN, which achieved an accuracy of 86.02%. Both Xpikeformer and its fully digital counterpart demonstrate increasing accuracy with longer encoding lengths. Xpikeformer shows converging accuracy (84.74%) at $T=10$, which is only 1% lower than the 85.68% accuracy achieved by SNN (Digital) at the same spike encoding length. The error bars indicate standard deviations from multiple tests, decreasing with longer encoding lengths for both SNN models. The uncertainty in the SNN (Digital) results arises from the Bernoulli encoder used in SSA. Xpikeformer exhibits higher uncertainty in accuracy performance compared to SNN (Digital) due to additional AIMC hardware noise, but this uncertainty reduces to less than 0.003% at $T\geq 10$, sufficiently small for stable inference.
Similarly, in Task 2, we observe that Xpikeformer shows converging BER when the spike encoding lengths $T\geq 6$, as shown in Fig. 8. Compared to the 0.051 BER achieved by the baseline ANN (Digital), SNN (Digital) and Xpikeformer achieve BERs of 0.061 and 0.067, respectively, at $T=6$, which are sufficiently low for the task. Increasing the spike coding length $T$ further improves accuracy slightly, with a tradeoff in latency and energy. This tradeoff should be considered in practical designs.
Figure 8: (a) CIFAR-10 classification accuracy in Task 1 and (b) bit error rate in Task 2 versus spike encoding length for SNN (Digital) and Xpikeformer. Baseline ANN results are shown for comparison.
VI-C Long-term Accuracy Performance
We evaluated the accuracy of memristor devices over time to investigate the impact of different training strategies and drift compensation methods. The experiment involved testing four spiking transformers, all based on the Xpikeformer architecture. These transformers were trained using two distinct strategies (CT and HWAT) and evaluated during inference with and without the application of the GDC compensation method. The results for Task 1 are shown in Fig. 9. Without GDC and HWAT, the accuracy started at 79.70% right after programming and showed a significant drop to below 70% within an hour. With HWAT but no GDC, the accuracy remained above 84% for at least one minute but decreased to below 80% within one hour due to conductance drift. With GDC but no HWAT, the model achieved up to 84.30% accuracy and maintained it above 70% within a year. With both GDC and HWAT, the accuracy was preserved at 81.13% after one year, dropping by only 3.6%. Similar results for Task 2 are shown in Fig. 9, demonstrating that the BER can be maintained at a low level, less than 0.08, for at least one year after writing weights to the devices when HWAT and GDC are applied. These findings suggest that with HWAT and GDC, Xpikeformer does not require frequent reprogramming to maintain high performance, thereby reducing maintenance costs and increasing the overall reliability of the hardware.
@@ VI-C Long-term Accuracy Performance
Figure 9: Accuracy performance of the hybrid implementation of SNN transformer trained with different strategies and drift compensation methods: conventional training (CT) + no compensation (NC), CT + global drift compensation (GDC), hardware-aware training (HWAT) + NC, and HWAT+GDC for (a) Task 1 and (b) Task 2.
@@ VII-A1 Baselines
We compare the runtime energy consumption of Xpikefomer with four baseline architectures: (i) ANN-Digital – a SOTA fully digital ANN transformer accelerator as reported in [31]; (ii) ANN-Hybrid – replacing the feed-forward and fully connected layers in (i) with AIMC crossbars. (iii) SNN-Digital (no SSA) – a fully digital implementation of SNN transformer with dot-product attention implemented with integer adders. (iv) SNN-Digital (with SSA) – a fully digital implementation of SNN transformer with dot-product attention implemented with SSA.
@@ VII-A2 Evaluation Methods
For the AIMC crossbars used in both Xpikeformer and the baseline configuration (ii), we customize the DNN+NeuroSim V1.4 framework [29] with the parameters specified in Table II to simulate the hardware implementations on a layer-by-layer basis. The area and energy estimates of the SSA tile were obtained by performing synthesis using the Cadence 45nm process design kit (PDK).
To estimate the baseline energy consumptions, we consider all digital operations (MAC, AC, MUL, ADD, and logic gates) using energy metrics based on 45 nm CMOS technology as reported in [51, 52], following the approach in [53]. We assume that there is sufficient on-chip cache to keep the model weights loaded for the entire duration of the hardware operation. Therefore, the energy consumption for loading model parameters is excluded from our runtime memory access energy calculations. For all memory access operations required to read or write input data, output data, and intermediate results, we assume that all necessary data is read from or written to the on-chip SRAM. For all input data, output data, and intermediate results in ANNs, as well as pre-activations in SNNs, we assume they are quantized to INT8 and stored in 8-bit SRAM when necessary. For the fully digital configurations (i), (iii), and (iv), we assume all model parameters are also quantized to INT8 and pre-loaded into the cache.
@@ VII-A3 Results
Figure 10: Energy consumption comparison between Xpikeformer and baseline implementations.
We evaluate the energy consumption on both Task 1 and Task 2. One main difference between Task 1 and Task 2 is the architectural parameters of the transformer networks, as detailed in Table I. For Task 1, a ViT-small encoder-only network is used, featuring 512 embedding dimensions, 6 layers, and 8 attention heads in each MHSA block. Task 2 employs a much smaller decoder-only network, differing from Task 1 by having 2 fewer layers, and a 2$\times$ lower embedding dimension while maintaining the same number of attention heads. Additionally, Task 2 employs a shorter spike encoding length during inference, $T=6$, compared to $T=10$ in Task 1, as this was determined to be the minimum length required to achieve the desired accuracy, as discussed in Section VI.
The energy consumption comparisons between Xpikeformer and the baseline implementations for Task 1 and Task 2 are shown in Fig.10 and Fig.10, respectively. Several key insights can be drawn from comparing the baseline implementations. Firstly, the spike encoding length plays a crucial role in converting ANN to SNN in the digital domain, with both computing energy and runtime memory access energy of a digital spiking transformer being directly proportional to it. In an ANN-Digital configuration, MAC operations account for over 99% of the total computing energy. By replacing these energy-intensive MAC operations with more efficient AC operations, the SNN-Digital (no SSA) achieves significantly lower computing energy – about 6.3$\times$ lower in Task 1 and 10.1$\times$ lower in Task 2. This reduction is approximately inversely proportional to the spike encoding length used in the tasks.
From a data communication perspective, while the softmax-free design of spiking transformers reduces some storage overhead for intermediate results, SNN-Digital (no SSA) incurs even higher runtime memory access energy (1.1$\times$ that of ANN-Digital) in Task 1. This is because the spike encoding length ($T=10$) adopted in Task 1 requires SNN-Digital (no SSA) to transfer binary data 10 times between layers, whereas the digital ANN only needs to transfer INT8 data once, making it less demanding in terms of memory access. Additionally, the burden of moving non-binary pre-activations 10 times significantly impacts SNN-Digital (no SSA), constituting approximately 40% of the total runtime memory access energy. Conversely, in Task 2, SNN-Digital (no SSA) consumes less runtime memory access energy (0.8$\times$ that of ANN-Digital) because intermediate results and pre-activations only need to be stored and retrieved $T=6$ times, compared to the $T=10$ cycles required in Task 1.
Considering the energy efficiency for computing, note that in Task 1, the attention operation in SNN-Digital (no SSA) consumes approximately 5 $\mu$J out of the total 43.1 $\mu$J computing energy. With the implementation of SSA, this 5 $\mu$J is significantly reduced to around 1 $\mu$J, resulting in a total computing energy of 38.4 $\mu$J. Similarly, in Task 2, attention contributes about 0.5 $\mu$J to the total computing energy in SNN-Digital (no SSA), which is reduced by 0.33 $\mu$J with SSA. Overall, SSA achieves a 3–5$\times$ reduction in energy consumption compared to spiking attention using AC operations.
The energy efficiency of SSA also stands out in terms of runtime memory access energy owing to the streamlined design of SSA tiles, which eliminates the need to store attention score $S^{t}$ and its pre-activations. This reduces the overall runtime memory access by 2$\times$ in Task 1 and 3.4$\times$ in Task 2.
To further improve the overall energy budget, the proposed Xpikeformer architecture incorporates an AIMC engine in addition to being equipped with SSA. This further reduces the overhead of intensive CMOS AC operations through AIMC, achieving the lowest computing energy consumption in both tasks. For runtime memory access energy, the streamlined integration between SAs and LIF units in Xpikeformer minimizes the need to write and read non-binary pre-activations over $T$ cycles, resulting in significantly reduced runtime memory access energy. Compared to SNN-Digital (SSA), Xpikeformer archives approximately 5.6$\times$ and 7.4$\times$ reduction in runtime memory access energy in Task 1 and Task 2, respectively.
Xpikeformer outperforms the other AIMC-integrated architecture, ANN-Hybrid, in energy efficiency primarily by eliminating the need for intermediate results storage. While ANN-Hybrid utilizes AIMC for feedforward operations and $Q$, $K$, and $V$ mapping, it still requires the storage of attention intermediate results ($QK^{\mathsf{T}}$, scaling results, softmax intermediate results, attention scores) and the output of the first layer in the feedforward blocks. These storage requirements are identical to those of ANN-Digital. In contrast, Xpikeformer eliminates the storage overhead of attention intermediate results by computing in the spike domain, aided by streamlined SSA tiles, and removes the need to store pre-activations in feedforward blocks through the integration of SAs and LIF units within the AIMC engine.
@@
It is worth noting that ANN-Hybrid exhibits competitive computing energy consumption. In Task 1, the computing energy of ANN-Hybrid (12.1 $\mu$J), which uses INT8 activations, is comparable to that of Xpikeformer (11.2 $\mu$J), which employs $T=10$. The slight overhead in ANN-Hybrid is due to the complexity of GeLu, normalization, and softmax operations. However, this advantage diminishes when the corresponding spiking transformer requires a shorter spike encoding length (compared to the ANN data quantization bitwidth) to achieve satisfactory accuracy. For example, in Task 2, Xpikeformer only needs $T=6$, while ANN-Hybrid requires 8 AIMC cycles to process INT8 feedforward operations. Additionally, the complex digital activation functions and normalization in ANN-Hybrid reduce its competitiveness in computing energy for this task.

### v2

@@ IV-C Dataflow
The AIMC engine serially drives input data through each forward propagation or fully connected layer. Each AIMC cycle, except for the first and last, includes two feedforward layers and one fully connected layer to project embeddings into the $QKV$ feature space. In each layer, data is processed in a token-wise event-driven manner: the spike encoding of each token embedding is input sequentially over $T$ time steps, followed by cycling through $N$ tokens. This approach allows the membrane potential stored in the LIF unit’s register to accumulate directly and reset upon token switching. Looping $N$ tokens for each time step $t$ would require storing and reading back membrane potentials, leading to unnecessary data transmission overhead.
@@
The SSA engine operates multiple SSA tiles in parallel, with each tile handling one head of the MHSA mechanism. As previously discussed, queries, keys, and values in SSA tiles are transmitted in a matrix-wise event-driven manner. Specifically, the three matrices $Q^{t}$, $K^{t}$, and $V^{t}$ are streamed to the SSA tile column by column over $d_{K}$ clock cycles. The subsequent matrices $Q^{t+1}$, $K^{t+1}$, and $V^{k+1}$ are transmitted in the next $d_{K}$ clock cycles. When pipelining in time steps, the computational delay of an SSA tile (from first input to first output) is approximately $d_{K}$ clock cycles.
@@ V-A Integrated Framework for Hybrid Training
We adopt a two-stage training procedure:
  1. 1.
    Conventional training (CT) is first performed in an ideal full-precision environment.
  2. 2.
    The trained parameters are then fine-tuned using hardware-aware training (HWAT), where noise is injected during forward propagation to simulate PCM-based hardware imperfections, while backward propagation remains ideal.
During training, forward propagation outputs are averaged over time steps to compute the loss, which follows a definition similar to that used in ANN transformer training. The network is trained using the AdamW optimizer [52].
@@ V-B Global Drift Compensation
To mitigate conductance drift in PCM devices within the AIMC engine, we implement a global drift compensation (GDC) method based on [53]. During the calibration phase, Xpikeformer periodically measures the output current of several SA columns using a known calibration input voltage. This allows Xpikeformer to track global shifts in conductance. During inference, the outputs are scaled by a factor derived from these calibration measurements. Calibration can be performed while tiles are idle, with current summation implemented within the system’s control unit to minimize overhead.
@@ VI Accuracy Evaluation
The inference accuracy of Xpikeformer is validated in hardware-simulated environments using AIHWKit. We evaluate the encoder-only and decoder-only architectures separately, each through a dedicated task.
VI-A Accuracy Results
Task 1 (Image Classification): To evaluate the accuracy performance of the encoder-only transformer, we select the image classification task, as it requires the model to extract and process relevant features from the entire input image in parallel. We train three different-sized vision transformers (ViTs) using the Xpikeformer implementation from scratch on the CIFAR10 and ImageNet-1K datasets, following the training strategy introduced in Section V-A. Inference accuracy is tested on the evaluation dataset using the simulation platform described in Section V-B. We compare the Xpikeformer results against two baseline implementations: (i) ANN-ViT (GPU) – Vanilla ViT trained with mixed precision (FP16/32) on GPU [2] and (ii) SNN-ViT (GPU) – SOTA software design of SNN-based ViT trained with FP16/32 on GPU [13]. The parameters of both baselines are INT8-quantized during the test stage.
The accuracy results are summarized in Table III. For SNN-based models, the minimum spike encoding length $T$ required for convergence (defined as $\Delta\text{Acc}<0.1$) is indicated in brackets. Xpikeformer generally requires a longer spike encoding length to converge compared to the SNN-ViT (GPU) due to hardware noise and stochasticity introduced by Bernoulli neurons. Despite this, Xpikeformer achieves competitive results, attaining 84.74% on CIFAR-10 and 70.50% on ImageNet, with accuracy gaps of less than 1.5% compared to same-sized ANN-ViT models.
While Xpikeformer consistently requires a longer spike encoding length than SNN-ViT (GPU) for the same task, the additional encoding overhead is smaller on ImageNet compared to CIFAR-10. This is because transformers generalize poorly on limited data but exhibit greater robustness to hardware noise and stochasticity when trained on larger, more diverse datasets.
Scaling up the model size further improves Xpikeformer’s performance while reducing the required spike encoding length. For instance, increasing the model size from 6-512 to 8-768 reduces the spike encoding length requirement from 8 to 7, while the accuracy gap relative to ANN-ViT decreases from 2.22% to 1.14%. This improvement occurs because the larger model dimensions effectively average out hardware noise and Bernoulli coding uncertainties.
TABLE III: Accuracy Performance on Image Classification Tasks
Model Size / Depth-Dim CIFAR10 ImageNet-1K
Accuracy ($T$) Accuracy ($T$)
ANN-ViT[2] (GPU) 4-384 83.41 -
6-512 86.02 65.58
8-768 - 71.64
SNN-ViT[13] (GPU) 4-384 82.83 (5) -
6-512 85.68 (4) 65.42 (6)
8-768 - 71.13 (4)
Xpikeformer-ViT (Simulated ASIC) 4-384 82.66 (11) -
6-512 84.74 (10) 63.36 (8)
8-768 - 70.50 (7)
TABLE IV: Accuracy Performance on In-Context Learning for Wireless Symbol Detection
Model Size / Depth-Dim $2\times 2$ Antennas $4\times 4$ Antennas
BER ($T$) BER ($T$)
ANN-GPT[3] (GPU) 4-256 0.051 0.141
8-512 0.049 0.077
SNN-GPT[30] (GPU) 4-256 0.061 (6) 0.196 (7)
8-512 0.058 (4) 0.086 (4)
Xpikeformer-GPT (Simulated ASIC) 4-256 0.067 (6) 0.205 (11)
8-512 0.063 (4) 0.091(5)
Task 2 (In-Context Learning for Wireless Communication): Decoder-only transformers are well-suited for tasks involving temporal sequential dependencies, particularly demonstrating the ability of in-context learning (ICL). In ICL, the model is presented with a sequence of query-answer pairs following a consistent pattern. Without explicit training on this pattern, it predicts the answer to a new query by attending to preceding query-answer pairs within the input sequence. This capability has been shown to enable wireless receivers to accurately classify symbols, even when the signals have passed through unknown wireless channels and been corrupted by nonlinearities [3, 4].
Following the formulation in [30], we adopt the two-symbol detection task, assuming 2$\times$2 and 4$\times$4 antenna configurations. A larger number of antennas implies higher computational complexity. Performance is evaluated using the bit error rate (BER) metric, where a lower BER indicates higher detection accuracy. The number of input query-answer pairs is fixed at 18. We compare the performance of Xpikeformer with the following baselines: (i) ANN-GPT (GPU) – a GPT-2-like decoder-only transformer running on GPUs [3], and (ii) SNN-GPT (GPU) – the SOTA software design of an SNN-based solution running on GPUs [30]. The baseline models are trained using FP16/32 precision and evaluated with INT8 quantization during the test phase.
The results are summarized in Table IV. On the 2$\times$2 antenna task, both 4-256 and 8-512 Xpikeformers achieve a low BER using the same spike encoding length ($T=6$ and $T=4$, respectively) as SNN-GPT (GPU). This is because the task, characterized by a small number of classes, is relatively simple. However, as the number of antennas increases, the number of classes grows exponentially, leading to degraded performance for both small-sized SNN-GPT and Xpikeformer-GPT models. Scaling up Xpikeformer-GPT from 4-256 to 8-512 brings greater performance gains than SNN-GPT, reducing BER by 0.114 with 6 fewer time steps, compared to a 0.110 reduction with 3 fewer time steps for SNN-GPT. This improvement is attributed to the larger model’s ability to better learn and average out hardware noise and stochastic variations.
VI-B Ablation Study: Long-term Accuracy Performance
@@
Figure 7: Long-term ImageNet accuracy of Xpikeformers trained with different strategies and drift compensation methods: conventional training (CT) + no compensation (NC), CT + global drift compensation (GDC), hardware-aware training (HWAT) + NC, and HWAT+GDC.
To evaluate the impact of NVM conductance drift on performance, we analyzed Xpikeformer’s accuracy over time under different training strategies and drift compensation methods. Four Xpikeformer-ViTs of size 8-768 were trained using all combinations of two training strategies—CT and HWAT—and two calibration settings—with and without GDC.
Without GDC or HWAT, accuracy started at 67.6% but dropped significantly to below 50% within one hour. With HWAT but no GDC, accuracy remained above 70% initially but rapidly declined to below 65% within an hour due to conductance drift. With GDC alone, accuracy improved to 69.9% and remained above 60% for one year. Combining HWAT and GDC achieved the best stability, with accuracy dropping by only 2.09% after one year. Similar trends were observed for other tasks tested in Section VI-A.
We also compared the long-term accuracy of Xpikeformer-ViTs (6-512 vs. 8-768), as shown in Table V. Despite having fewer parameters, the 6-512 model exhibits a larger accuracy drop (--2.95% after one year) compared to the 8-768 model. The smaller model also gained less benefit from GDC alone. This is because larger model dimensions help average out hardware noise uncertainties and the stochastic component of conductance drift during GDC calibration.
TABLE V: Xpikeformer Long-Term (One Year) Accuracy on ImageNet-1K
Size CT+NC HWAT+NC CT+GDC HWAT+GDC
6-512 9.63 (-56.73) 11.32 (-55.04) 54.37 (-11.99) 60.41 (-2.95)
8-768 6.10 (-64.40) 5.46 (-65.04) 62.80 (-7.70) 68.41 (-2.09)
@@ VII-A1 Baselines
We compare the runtime energy consumption of Xpikeformer with three baseline architectures: (i) ANN-Quant – a SOTA fully digital accelerator for ANN-based transformer models, as reported in [34]. This baseline serves as a benchmark for fully digital ANN architectures. (ii) ANN-Quant+AIMC – this baseline modifies (i) by replacing the feed-forward and fully connected layers with AIMC crossbars based on PCM, reducing energy costs associated with MAC operations. (iii) SNN-Digi-Opt – this baseline assumes an ideal digital implementation of spiking transformer using the arithmetic operations proposed in [15].
@@ VII-A2 Evaluation Methods
For the AIMC crossbars used in both Xpikeformer and the ANN-Quant+AIMC, we customize the DNN+NeuroSim V1.4 framework [32] with the parameters specified in Table II to simulate the hardware implementations on a layer-by-layer basis. The area and energy estimates of the SSA tile were obtained by performing synthesis using the Cadence 45nm process design kit (PDK).
To estimate the baseline energy consumptions, we consider all digital operations (MAC, AC, MUL, ADD, and logic gates) using energy metrics based on 45 nm CMOS technology as reported in [54, 55], following the approach in [56]. We assume that there is sufficient on-chip cache to keep the model weights loaded for the entire duration of the hardware operation. Therefore, the energy consumption for loading model parameters is excluded from our runtime memory access energy calculations. For all memory access operations required to read or write input data, output data, and intermediate results, we assume that all necessary data is read from or written to the on-chip SRAM. For all input data, output data, and intermediate results in ANNs, as well as pre-activations in SNNs, we assume they are quantized to INT8 and stored in 8-bit SRAM when necessary. For the fully digital configurations ANN-Quant and SNN-Digi-Opt, we assume all model parameters are also quantized to INT8 and pre-loaded into the cache.
For SNN implementations, as both computing energy and runtime memory access energy are directly proportional to the spike encoding length $T$, we use the minimum required spike encoding length $T$ for Xpikeformer and SNN-Digi-Opt, respectively, as determined in Section VI-A, to achieve their converged accuracy. This ensures a fair comparison of computational efficiency, as both per-inference energy consumption and latency scale directly with the spike encoding length in Xpikeformer and the baseline hardware implementations considered.
@@ VII-A3 Results
Figure 8: Energy consumption comparison between Xpikeformer and baseline implementations on downstream application (a) ImageNet-1K and (b) ICL symbol detection ($4\times 4$ antennas).
We evaluate the energy consumption on both Task 1 (ImageNet) and Task 2 (4$\times$4 antennas) across different model sizes. The energy consumption comparisons between Xpikeformer and the baseline implementations are shown in Fig. 8. It can be observed that Xpikeformer consistently consumes less energy than all three baseline models across varying model sizes. Specifically, on ImageNet-1K, Xpikeformer achieves 9.6-13$\times$ energy reduction than ANN-Quant and 5.4-5.9$\times$ energy reduction than ANN-Quant+AIMC. Notably, despite using more timesteps to achieve comparable accuracy, Xpikeformer still consumes 1.8-1.9$\times$ less energy than SNN-Digi-Opt. Similar results go for the ICL symbol equalization task, where Xpikeformer achieves energy reduction 7-14.5$\times$ than ANN-Quant, 2.6-4.0$\times$ than ANN-Quant+AIMC, and 1.6-1.8$\times$ than SNN-Digi-Opt. As the model scales up, the energy efficiency becomes more evident due to the reduced requirement for spike encoding length $T$.
Several key insights can be drawn from comparing computational energy consumption. Firstly, the results for ANN-Quant demonstrate that MAC operations are highly energy-hungry, usually dominating more than 90% computing energy even when operating at low INT8 precision. This energy overhead represents a bottleneck that traditional quantized ANN-based transformer accelerators cannot overcome. In contrast, the other implementations address this challenge using different approaches: ANN-Quant+AIMC and Xpikeformer leverage AIMC to achieve efficient MVM in fully connected and linear layers, while SNN-Digi-Opt replaces MAC operations with masked ADD operations, taking advantage of the binary nature of SNN activations. Additionally, both SNN implementations avoid energy-intensive operations such as softmax and layer normalization, which remain necessary in ANN accelerators.
Although Xpikeformer requires a few additional timesteps, it still consumes less computational energy than SNN-Digi-Opt. This advantage stems from two key factors: (1) For the attention mechanism, SNN-Digi-Opt uses masked addition and integer multiplication to compute attention scores, whereas Xpikeformer employs a more energy-efficient logic-gate-based SSA engine. (2) For feed-forward layers, while AIMC requires complex peripheral circuitry and ADCs, but is generally more energy-efficient per timestep than the digital masked addition approach. Specifically, AIMC computes MVM with $O(1)$ complexity, whereas the digital method requires $O(n)$ operations, with $n$ being the dimension of the vector.
@@
From the runtime memory access perspective, we observe that both ANN-Quant and ANN-Quant+AIMC consume the same high amount of memory access energy, as AIMC does not reduce intermediate data storage overhead. In contrast, both SNN implementations achieve lower memory access energy compared to ANN accelerators, primarily owing to the softmax-free SNN attention mechanism and the binary nature of SNN activations. However, this memory efficiency benefit diminishes when the number of timesteps $T$ exceeds the bit width of the ANN activation datatype.
Xpikeformer achieves significantly lower memory access energy compared to SNN-Digi-Opt, because (1) SNN-Digi-Opt requires storing $T$ non-binary pre-activations from every crossbar output column before performing LIF calculations, whereas Xpikeformer’s rowblock-wise mapping strategy in its AIMC engine eliminates this bottleneck. (2) The streamlined design of the SSA engine eliminates the need to store intermediate attention scores.